A 50.00 mL sample of gas at 15.0℃ is heated to 35.0℃ at constant pressure. What is the new volume of the gas in mL?
Answer #1

This is a simple example of Charles' law

V1/T1=V2/T2. V1 and T1 are initial values and V2 and T2 are final values.

which states that the volume of a gas is directly proportional to its temperature in K.

Here V1=50ml. ,T1=15℃=15+273.15=288.15K

T2=35℃=35+273.15= 308.15K .We have to find V2

From the equation V2=(V1×T2)/T1

=(50×308.15)/288.15 = 53.47ml

New volume of gas = 53.47ml